OpenStudy (anonymous):

\[\\int \frac{1-2cos x}{sin^2x}dx=\] How to get out of the fraction?

OpenStudy (amistre64):

i might consider multiplying by 2/2

OpenStudy (amistre64):

or is that a different identity that im thinking of

OpenStudy (anonymous):

divide each term by \[\sin^2(x)\] would do it

OpenStudy (amistre64):

1/sin^2 = csc^2 which ints up by basic trig stuff -2cos/sin^2 tho ... is 2 cot(x)csc(x)

OpenStudy (amistre64):

yes, it would do it :)

OpenStudy (anonymous):

good morning lots of new features i see left side looks like a three ring circus

OpenStudy (amistre64):

yeah, its a bit of a distraction to me

OpenStudy (amistre64):

also, to police or find questions now i have to flip back and forth between open and closed

OpenStudy (eyust707):

yes but now people dont post a million ?s at once

OpenStudy (anonymous):

\[\int\frac{1}{\sin^2(x)}dx-2\int \frac{\cos(x)}{\sin^2(x)}dx\]

OpenStudy (amistre64):

they still post them, they just have them all lined up in the closed list

OpenStudy (anonymous):

it would be nice if there was an explanation of the new features, it is hard to teach an old dog new tricks

OpenStudy (amistre64):

:) theres a new release posted in the feedback section

OpenStudy (eyust707):

=P

OpenStudy (anonymous):

\[\int\frac{1}{\sin^2(x)}dx=\int csc^2(x)dx=-\cot(x) \]second one is a u-sub

OpenStudy (anonymous):

http://mycalculus.com/calculus_solver.aspx